# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v6

@@ C.1 Attraction parameter for Outside Option: $v_{0}=\Theta(1)$ is Common in Contextual MNL Bandits
Then, by dividing the denominator and numerator by $\exp\left((x^{\prime}_{t0})^{\top}\mathbf{w}^{\star}\right)$, and defining $x_{ti}:=x^{\prime}_{ti}-x^{\prime}_{t0}$, we obtain the MNL probability in the form presented in (1) with $v_{0}=\exp(0)=1$. Note that this division does not change the probability. Therefore, $v_{0}=\Theta(1)$ is natural and common in contextual MNL bandit literature.
@@ C.2 Self-concordant-like Function
Definition C.1 (Self-concordant-like function, Tran-Dinh et al. 48).
@@
Then, the MNL loss defined in (2) is $3\sqrt{2}$-self-concordant-like function.
@@
For any $t\in[T]$, the multinomial logistic loss $\ell_{t}(\mathbf{w})$, defined in (2), is $3\sqrt{2}$-self-concordant-like.
@@
Plugging in (C.3) into (C.1), we obtain
@@
Now, we are ready to prove the proposition. For any $t\in[T]$, let $n=|S_{t}|$ and $c_{1}=x_{ti_{1}},c_{2}=x_{ti_{2}},\dots,c_{n}=x_{ti_{n}}$. Define a function $f\in\mathcal{C}^{3}:\mathbb{R}^{d}\rightarrow\ \mathbb{R}$ as $f(\bm{\theta}):=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}}\right)$. Let $\bm{\delta}\in\mathbb{R}^{d}$ and let $f(\bm{\theta}+s\bm{\delta})=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}+sc_{i}^{\top}\bm{\delta}}\right)=\log\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)=\phi(s)$, where $a_{i}=c_{i}^{\top}\bm{\delta}$, $b_{i}=c_{i}^{\top}\bm{\theta}$ for $i=1,\dots,n$, and $a_{i}=0$ and $b_{i}=\log v_{0}$ for $i=0$. Then, by (C.4), we get
@@
where the last inequality holds due to Assumption  1 that $\|c_{i}\|_{2}=\|x_{tj_{i}}\|_{2}\leq 1$. Then, by Definition C.1, $f$ is $3\sqrt{2}$-self-concordant-like. Since $\ell_{t}$ is the sum of $f$ and a linear operator, which has third derivatives equal to zero, it follows that $\ell_{t}$ is also $3\sqrt{2}$-self-concordant-like function. ∎
@@
Lemma C.1 (Theorem 3 of Tran-Dinh et al. 48).
@@ D.1 Adversarial Construction and Bayes Risk
Let $\epsilon\in(0,1/d\sqrt{d})$ be a small positive parameter to be specified later. For every subset $V\subseteq[d]$, we define the corresponding parameter $\mathbf{w}_{V}\in\mathbb{R}^{d}$ as $[\mathbf{w}_{V}]_{j}=\epsilon$ for all $j\in V$, and $[\mathbf{w}_{V}]_{j}=0$ for all $j\notin V$. Then, we consider the following parameter set
@@
The context vectors $\{x_{ti}\}$ are constructed to be invariant across rounds $t$. For each $t$ and $U\in\mathcal{V}_{d/4}$, $K$ identical context vectors  [^4] Recall that $K$ is the maximum allowed assortment capacity. $x_{U}$ are constructed as follows:
@@
For all $V,U\in\mathcal{V}_{d/4}$, it can be verified that $\mathbf{w}_{V}$ and $x_{U}$ satisfy the requirements of a bounded assumption 1 as follows:
@@ D.2 Main Proof of Theorem 1
$\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (D.2)
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}$ $\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$
@@
Now we bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in (D.3) for any $V\in\mathcal{V}_{d/4-1}$. For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$ denote the laws under $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup j}$, respectively. Then, we have
@@
Therefore, combining (D.3), (D.4), and Lemma D.2, we have
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}$ $\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$
@@ D.3 Proofs of Lemmas for Theorem 1
where the inequality holds since $\max\left\{\exp\left(x^{\top}\mathbf{w}_{V}\right),\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right\}\leq e$. To further bound the right-hand side of  (D.5), we use the fact that $21+a\leq e^{a}\leq 1+a+a^{2}/$ 2 for all $a\in[0,1]$, which can be easily shown by Taylor expansion. Thus, we get
@@
where the first inequality holds because $\log(1+x)\leq x$ for all $x>-1$.
@@
Combining (D.6), (D.7), and (D.8), we derive that
@@ E.1 Main Proof of Theorem 2
where the inequality is by Lemma E.3. To bound the first term in (E.3), by applying the AM-GM inequality, we get
@@
By plugging (E.4) into (E.3), we have
@@
Now, we bound each term of (E.6) respectively. For the first term, we decompose it as follows:
@@
To bound the first term on the right-hand side of (E.7), we apply the Cauchy-Schwarz inequality.
@@
Now, we bound the second term on the right-hand side of (E.7). Let the virtual context for the outside option be $x_{t0}=\mathbf{0}$. Then, by the mean value theorem, there exists $\bm{\xi}_{t}=(1-c)\mathbf{w}^{\star}+c\mathbf{w}_{t+1}$ for some $c\in(0,1)$ such that
@@
Finally, we bound the third term on the right-hand side of (E.7). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
where the last inequality holds by Lemma E.2. By plugging (E.8), (E.9), and (E.10) into (E.7) and multiplying $2\beta_{T}(\delta)$, we get
@@
Moreover, by applying Lemma E.2, we can directly bound the second term of (E.6).
@@
Finally, plugging (E.11) and (E.12) into (E.6), we obtain
@@
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, since $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K\right)$, we have $\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(Be^{B}\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T}+\frac{1}{\kappa}d^{2}\right).$ It’s important to note that one of our main goals is to explicitly demonstrate the regret depends on $K$ and $v_{0}$. In deriving such a result, the dependence on $e^{B}$ is unavoidable to our best knowledge. Note that for non-uniform rewards, the regret bound does not depend on $e^{B}$ (refer Remark H.1).
@@ E.2 Proofs of Lemmas for Theorem 2
To establish the proof for the second inequality, we return to (E.13):
@@
Let $p_{i}(\mathbf{u})=\frac{e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ and $p_{0}(\mathbf{u})=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ For $i=j$, we have
@@ F.1 Main Proof of Lemma 1
where ~$\tilde{\ell}_{t}(\mathbf{w})=\ell_{t}(\mathbf{w}_{t})+\langle\mathbf{w}-\mathbf{w}_{t},\nabla\ell_{t}(\mathbf{w}_{t})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{t}\|^{2}_{\nabla^{2}\ell_{t}(\mathbf{w}_{t})}$ and $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Let $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda>0$. Then, we have
@@
$-\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}.$ (F.1)
We first bound the first term in (F.1). For simplicity, we define the softmax function at round $t$ $\bm{\sigma}_{t}(\mathbf{z}):\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$ as follows:
@@
where $[\cdot]_{i}$ denotes $i$’th element of the the input vector. We denote the probability of choosing the outside option as $[\bm{\sigma}_{t}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})}$. Although $[\bm{\sigma}_{t}(\mathbf{z})]_{0}$ is not the output of the softmax function $\bm{\sigma}_{t}(\mathbf{z})$, we represent it in a form similar to that in (F.2) for simplicity. Then, the user choice model in (1) can be equivalently expressed as $p_{t}(i|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{i}$ for all $i\in[K_{t}]$ and $p_{t}(0|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{0}$. Furthermore, the loss function (2) can also be written as $\ell(\mathbf{z}_{t},\mathbf{y}_{t})=\sum_{k=0}^{K_{t}}\mathbf{1}\left\{y_{ti}=1\right\}\cdot\log\left(\frac{1}{[\bm{\sigma}_{t}(\mathbf{z}_{t})]_{k}}\right)$.
Define a pseudo-inverse function of $\bm{\sigma}_{t}(\cdot)$ as $\bm{\sigma}_{t}^{+}:\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$, where $[\bm{\sigma}_{t}^{+}(\mathbf{q})]_{i}=\log\left(q_{i}/(1-\|\mathbf{q}\|_{1})\right)$ for any $\mathbf{q}\in\{\mathbf{p}\in[0,1]^{K_{t}}\mid\|\mathbf{p}\|_{1}<1\}$. Then, inspired by the previous studies on binary logistic bandit [21], we decompose the regret into two terms by introducing an intermediate term.
@@
where $\tilde{\mathbf{z}}_{s}:=\bm{\sigma}_{s}^{+}\left(\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[\bm{\sigma}_{s}\left((x_{sj}^{\top}\mathbf{w})_{j\in S_{s}}\right)\right]\right)$, and $P_{s}:=\mathcal{N}(\mathbf{w}_{s},(1+cH_{s}^{-1}))$ is the Gaussian distribution with mean $\mathbf{w}_{s}$ and covariance matrix $cH_{s}^{-1}$, where $c>0$ is a positive constant to be specified later. We first show that the term $(a)$ is bounded by $\mathcal{O}\left(\log K\log t\right)$ with high probability.
@@
Let $\delta\in(0,1]$ and $\lambda\geq 1$. Under Assumptions 1, for all $t\in[T]$, with probability at least $1-\delta$, we have
$\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})$
$\leq\left(3\log(1+(K+1)t)+3\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+2.$
@@
$\leq 2\eta\Bigg{[}\!\left(3\log(1+(K+1)t)+3\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)$
$+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg{]}+\!4\lambda\!+12\sqrt{2}\eta\sum_{s=1}^{t}\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}+\left(\frac{\eta}{c}\!-\!1\right)\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$\leq 2\eta\Bigg{[}\!\left(3\log(1+(K+1)t)+3\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)$
$+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg{]}+\!4\lambda=:\beta_{t+1}(\delta)^{2},$
@@
$12\sqrt{2}\eta\sum_{s=1}^{t}$ $\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}+\left(\frac{\eta}{c}\!-\!1\right)\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$=12\sqrt{2}\eta\sum_{s=1}^{t}\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}-\frac{1}{7}\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$

### v9

@@ C.1 Attraction parameter for Outside Option: $v_{0}=\Theta(1)$ is Common in Contextual MNL Bandits
Then, by dividing the denominator and numerator by $\exp\left((x^{\prime}_{t0})^{\top}\mathbf{w}^{\star}\right)$, and defining $x_{ti}:=x^{\prime}_{ti}-x^{\prime}_{t0}$, we obtain the MNL probability in the form presented in Equation (1) with $v_{0}=\exp(0)=1$. Note that this division does not change the probability. Therefore, $v_{0}=\Theta(1)$ is natural and common in contextual MNL bandit literature.
@@ C.2 Self-concordant-like Function
Definition C.1 (Self-concordant-like function, Tran-Dinh et al. 47).
@@
Then, the MNL loss defined in Equation (2) is $3\sqrt{2}$-self-concordant-like function.
@@
For any $t\in[T]$, the multinomial logistic loss $\ell_{t}(\mathbf{w})$, defined in Equation (2), is $3\sqrt{2}$-self-concordant-like.
@@
Plugging in Equation (C.3) into Equation (C.1), we obtain
@@
Now, we are ready to prove the proposition. For any $t\in[T]$, let $n=|S_{t}|$ and $c_{1}=x_{ti_{1}},c_{2}=x_{ti_{2}},\dots,c_{n}=x_{ti_{n}}$. Define a function $f\in\mathcal{C}^{3}:\mathbb{R}^{d}\rightarrow\ \mathbb{R}$ as $f(\bm{\theta}):=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}}\right)$. Let $\bm{\delta}\in\mathbb{R}^{d}$ and let $f(\bm{\theta}+s\bm{\delta})=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}+sc_{i}^{\top}\bm{\delta}}\right)=\log\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)=\phi(s)$, where $a_{i}=c_{i}^{\top}\bm{\delta}$, $b_{i}=c_{i}^{\top}\bm{\theta}$ for $i=1,\dots,n$, and $a_{i}=0$ and $b_{i}=\log v_{0}$ for $i=0$. Then, by Equation (C.4), we get
@@
where the last inequality holds due to Assumption 1 that $\|c_{i}\|_{2}=\|x_{tj_{i}}\|_{2}\leq 1$. Then, by Definition C.1, $f$ is $3\sqrt{2}$-self-concordant-like. Since $\ell_{t}$ is the sum of $f$ and a linear operator, which has third derivatives equal to zero, it follows that $\ell_{t}$ is also $3\sqrt{2}$-self-concordant-like function. ∎
@@
Lemma C.1 (Theorem 3 of Tran-Dinh et al. 47).
@@ D.1 Adversarial Construction and Bayes Risk
Let $\epsilon\in(0,1/d\sqrt{d})$ be a small positive parameter to be specified later. For every subset $V\subseteq[d]$, we define the corresponding parameter $\mathbf{w}_{V}\in\mathbb{R}^{d}$ as $\mathbf{w}_{V}]_{j}=\epsilon$ for all $j\in V$, and $\mathbf{w}_{V}]_{j}=0$ for all $j\notin V$. Then, we consider the following parameter set
@@
The context vectors $x_{ti}$ are constructed to be invariant across rounds $t$. For each $t$ and $U\in\mathcal{V}_{d/4}$, $K$ identical context vectors  [^4] Recall that $K$ is the maximum allowed assortment capacity. $x_{U}$ are constructed as follows:
@@
For all $V,U\in\mathcal{V}_{d/4}$, it can be verified that $\mathbf{w}_{V}$ and $x_{U}$ satisfy the requirements of a bounded assumption (Assumption 1) as follows:
@@ D.2 Main Proof of Theorem 1
$\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (D.2)
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}$ $\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
@@
Now we bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in Equation (D.3) for any $V\in\mathcal{V}_{d/4-1}$. For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$ denote the laws under $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup j}$, respectively. Then, we have
@@
Therefore, combining Equation (D.3), Equation (D.4), and Lemma D.2, we have
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}$ $\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
@@ D.3 Proofs of Lemmas for Theorem 1
where the inequality holds since $\max\left\{\exp\left(x^{\top}\mathbf{w}_{V}\right),\exp\left(\hat{x}^{\top}\mathbf{w}_{V}\right)\right\}\leq e$. To further bound the right-hand side of Equation (D.5), we use the fact that $1+a\leq e^{a}\leq 1+a+a^{2}/2$ for all $a\in[0,1]$, which can be easily shown by Taylor expansion. Thus, we get
@@
where the first inequality holds because $\log(1+x)\leq x$ for all x>1x>-1.
@@
Combining Equation (D.6), (D.7), and (D.8), we derive that
@@ E.1 Main Proof of Theorem 2
where the inequality is by Lemma E.3. To bound the first term in Equation (E.3), by applying the AM-GM inequality, we get
@@
By plugging Equation (E.4) into Equation (E.3), we have
@@
Now, we bound each term of Equation (E.6) respectively. For the first term, we decompose it as follows:
@@
To bound the first term on the right-hand side of Equation (E.7), we apply the Cauchy-Schwarz inequality.
@@
Now, we bound the second term on the right-hand side of Equation (E.7). Let the virtual context for the outside option be $x_{t0}=\mathbf{0}$. Then, by the mean value theorem, there exists $\bm{\xi}_{t}=(1-c)\mathbf{w}^{\star}+c\mathbf{w}_{t+1}$ for some $c\in(0,1)$ such that
@@
Finally, we bound the third term on the right-hand side of Equation (E.7). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
where the last inequality holds by Lemma E.2. By plugging Equation (E.8), (E.9), and (E.10) into Equation (E.7) and multiplying $2\beta_{T}(\delta)$, we get
@@
Moreover, by applying Lemma E.2, we can directly bound the second term of Equation (E.6).
@@
Finally, plugging Equation (E.11) and (E.12) into Equation (E.6), we obtain
@@
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, since $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right)$ (refer Corollary F.1), we have $\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(B^{3/2}e^{B}\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T}+\frac{1}{\kappa}B^{3}d^{2}\right).$ It’s important to note that one of our main goals is to explicitly demonstrate the regret depends on $K$ and $v_{0}$. In deriving such a result, the dependence on $e^{B}$ is unavoidable to our best knowledge. Note that for non-uniform rewards, the regret bound does not depend on $e^{B}$ (refer Remark H.1).
@@ E.2 Proofs of Lemmas for Theorem 2
To establish the proof for the second inequality, we return to Equation (E.13):
@@
Let $p_{i}(\mathbf{u})=\frac{e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ and $p_{0}(\mathbf{u})=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ . For $i=j$, we have
@@ F.1 Main Proof of Lemma 1
where $\tilde{\ell}_{t}(\mathbf{w})=\ell_{t}(\mathbf{w}_{t})+\langle\mathbf{w}-\mathbf{w}_{t},\nabla\ell_{t}(\mathbf{w}_{t})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{t}\|^{2}_{\nabla^{2}\ell_{t}(\mathbf{w}_{t})}$ and $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Let $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda>0$. Then, we have
@@
$-\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}.$ (F.1)
We first bound the first term in Equation (F.1). For simplicity, we define the softmax function at round $t$ $\bm{\sigma}_{t}(\mathbf{z}):\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$ as follows:
@@
where $\cdot]_{i}$ denotes $i$’th element of the input vector. We denote the probability of choosing the outside option as $\bm{\sigma}_{t}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})}$. Although $\bm{\sigma}_{t}(\mathbf{z})]_{0}$ is not the output of the softmax function $\bm{\sigma}_{t}(\mathbf{z})$, we represent it in a form similar to that in Equation (F.2) for simplicity. Then, the user choice model in Equation (1) can be equivalently expressed as $p_{t}(i|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{i}$ for all $i\in[K_{t}]$ and $p_{t}(0|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{0}$. Furthermore, the loss function in Equation (2) can also be written as $\ell(\mathbf{z}_{t},\mathbf{y}_{t})=\sum_{k=0}^{K_{t}}\mathbf{1}\left\{y_{ti}=1\right\}\cdot\log\left(\frac{1}{[\bm{\sigma}_{t}(\mathbf{z}_{t})]_{k}}\right)$.
Define a pseudo-inverse function of $\bm{\sigma}_{t}(\cdot)$ as $\bm{\sigma}_{t}^{+}:\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$, where $\bm{\sigma}_{t}^{+}(\mathbf{q})]_{i}=\log\left(q_{i}/(1-\|\mathbf{q}\|_{1})\right)$ for any $\mathbf{q}\in\{\mathbf{p}\in[0,1]^{K_{t}}\mid\|\mathbf{p}\|_{1}<1\}$. Then, inspired by the previous studies on binary logistic bandit [21], we decompose the regret into two terms by introducing an intermediate term.
@@
where $\tilde{\mathbf{z}}_{s}:=\bm{\sigma}_{s}^{+}\left(\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[\bm{\sigma}_{s}\left((x_{sj}^{\top}\mathbf{w})_{j\in S_{s}}\right)\right]\right)$, and $P_{s}:=\mathcal{N}(\mathbf{w}_{s},cH_{s}^{-1})$ is the Gaussian distribution with mean $\mathbf{w}_{s}$ and covariance matrix $cH_{s}^{-1}$, where $c>0$ is a positive constant to be specified later. We first show that the term $(a)$ is bounded by $\mathcal{O}\left(\log K(\log t)^{2}\right)$ with high probability.
@@
Let $\delta\in(0,1]$. Under Assumptions 1, for all $t\in[T]$, with probability at least $1-\delta$, we have
$\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})$ $\leq 11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2.$
@@
$\leq 2\eta\Bigg[\!11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg]$
$+\!4\lambda\!+12\sqrt{2}\eta\sum_{s=1}^{t}\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}+\left(\frac{\eta}{c}\!-\!1\right)\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$\leq 2\eta\Bigg[\!11\cdot\left(3\log(1+(K+1)t)+3\right)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2+\sqrt{6}cd\log\left(1+\frac{t+1}{2\lambda}\right)\!\Bigg]$
$+\!4\lambda\!=:\beta^{\prime}_{t+1}(\delta)^{2}=\mathcal{O}\left(d(\log t\log K)^{2}\right),$ (F.4)
@@
$12\sqrt{2}\eta\sum_{s=1}^{t}$ $\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}+\left(\frac{\eta}{c}\!-\!1\right)\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$
$=12\sqrt{2}\eta\sum_{s=1}^{t}\!\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}-\frac{1}{7}\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}$